Grade 12Algebra

Give first the step you will use to separate the variable and then solve the equation:
b/2 = 6

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Now we have to multiply both sides of the equation by 2,
Then we get,
= b/2 × 2= 6 × 2
= b = 12

Thanks